OpenStudy (anonymous):

Rationalize the denominator of Sqrt of -9/ (4-7i)-(6-6i)

OpenStudy (anonymous):

Please helpppp

OpenStudy (anonymous):

combine like terms in the denominator first

OpenStudy (anonymous):

-13i

OpenStudy (anonymous):

you should get \[\frac{3i}{-2-i}\] if i am reading it correctly

OpenStudy (anonymous):

No

OpenStudy (anonymous):

That is not in my answers

Nnesha (nnesha):

first you have to distribute bracket by negative sign

OpenStudy (anonymous):

i didn't say that was the answer, i said that was the first step there is more to go

OpenStudy (anonymous):

ok

Nnesha (nnesha):

because you did -7i -6i instead -7i + 6i

OpenStudy (anonymous):

-i

OpenStudy (anonymous):

is this the denominator \[(4-7i)-(6-6i) \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

remove the parentheses and combine like terms you should get what i wrote above try it

OpenStudy (anonymous):

then there is a couple more steps to write this in standard form

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I got that what do i do now

OpenStudy (anonymous):

\[\frac{3i}{-2-i}\] multiply top and bottom by the conjugate of the denominator, which is \(-2+i\)

OpenStudy (anonymous):

\[\frac{3i}{-2-i}\times \frac{-2+i}{-2+i}\]

OpenStudy (anonymous):

the denominator will be \(2^2+1^2=5\)

OpenStudy (anonymous):

the numerator will be whatever you get when you multiply

OpenStudy (anonymous):

So -3-6i/5

OpenStudy (anonymous):

yeah looks good

OpenStudy (anonymous):

Ca u help with another

OpenStudy (anonymous):

Which of the following is teh conjugate of a complex number with 2 as the real part and -4 as the imaginary part.

OpenStudy (anonymous):

I think its 2+4i